OpenStudy (anonymous):

y varies directly as the square of x. When x = 1, y = 8. Find x when y = 128. Answer x = ±8 x = ±16 x = ±4 x = ±6

OpenStudy (anonymous):

y=kx^2 k=8 x^2 = 128/8 = 16 \[x = \pm4\]